The magnetic field produced by the solenoid in a magnetic resonance imaging (MRI) system designed for measurements on whole human bodies has a field strength of 8.0 T, and the current in the solenoid is 1.5 102 A. What is the number of turns per meter of length of the solenoid? Note that the solenoid used to produce the magnetic field in this type of system has a length that is not very long compared to its diameter. Because of this and other design considerations, your answer will be only an approximation
Answer #1

Magnetic field inside solenoid is given by

B = u0*n*i

n = N/L = number of turns per meter of the length

n = B/(u0*i)

n = 8/(4*pi*10^-7*1.5*10^2)

n = 42441 Turns = 4.244*10^4 Turns/m